OpenStudy (anonymous):

find the required number TU=3 RS=15 TW=4 TR=9 WS=

OpenStudy (anonymous):

Heres a photo for referecnce

OpenStudy (kinggeorge):

You know that TU=3, and TR=9 Can you tell me what RU is?

OpenStudy (anonymous):

actually i can't RU is also anotheer question i was gonna ask. but can you help me find both

OpenStudy (kinggeorge):

Sure. Look at your triangle. Notice that TR=TU+RU. You're given that TR=9, and TU=3. Do the substitutions, you get \[9=3+RU\]Can you solve for RU from there?

OpenStudy (anonymous):

kind of

OpenStudy (kinggeorge):

Why don't you give it a shot.

OpenStudy (anonymous):

ok would it be 27i tried to cross multiply

OpenStudy (kinggeorge):

Not quite. You don't want to cross multiply anything yet. You only need to subtract 3 from both sides.\[9=3+RU\]\[9-3=3-3+RU\]\[6=0+RU\]\[RU=6\]

OpenStudy (kinggeorge):

Now that we have RU, it's much easier to find WS. You're given \[\frac{TU}{UR}=\frac{TW}{WS}\]You are also given TU=9, TW=4, and we just found UR=6. Do the substitutions, we get \[\frac{9}{3}=\frac{4}{WS}\]Can you solve from here? (Hint: cross multiply)

OpenStudy (anonymous):

i got 1.3 repeating

OpenStudy (kinggeorge):

That looks perfect. Good job!

OpenStudy (kinggeorge):

Wait!

OpenStudy (anonymous):

what???

OpenStudy (kinggeorge):

I did some things wrong in my substitution :( Completely my fault. Hold on one second to let me correct everything.

OpenStudy (kinggeorge):

You are also given TU=3, TW=4, and we just found UR=6. Do the substitutions, we get \[\frac{3}{6}=\frac{4}{WS}\]Now can you solve this (cross multiply again). Sorry about the confusion. That was completely my fault.

OpenStudy (anonymous):

its ok and thank you btw do you know anything about algerbra and ratios

OpenStudy (kinggeorge):

I can handle my algebra and ratios.

OpenStudy (anonymous):

use the figure to find each ratio in the simplest form

OpenStudy (kinggeorge):

First off, can you tell me what AB is? Hint: \[AB=AD+DB\]

OpenStudy (anonymous):

18

OpenStudy (kinggeorge):

Right on. Now we're given that AD=12, and we've found AB=18. We're told to calculate \[\frac{AD}{AB}=\frac{12}{18}\]What is this in simplest form?

OpenStudy (anonymous):

6/9

OpenStudy (kinggeorge):

It can be reduced a little bit farther than that.

OpenStudy (kinggeorge):

Sorry, but I've got to go now. If you have more questions close this one and ask to the left. One final hint: The denominator will be 3.